It takes a boat 3 hours to travel down a river from Point A to Point B and 5 hours to travel up the river from Point B to Point A. How long would it take the same boat to go from A to B in still water?
Answer by mananth(16946)

x speed of boat in still water
y speed of current
down river speed = x+y
upstream speed = x-y
Distance = same upstream and downstream
d=rt
3(x+y)=5(x-y)
3x+3y=5x-5y
2x=8y
x=4y
distance = 3(x+y)
but x=4y
distance = 3(4y+y)
=15y
speed in still water = 4y
time = d/r
=15y/4y
=3.75 hours
Time in still water = 3.75 hours